2-20 Scattergraph; high-low method; cost estimation (LO 2) The Aust Corporation has gathered the following data on its copy m please answer A - E! Thank you!
Answer #1

Saltorgraph & Equations my Choo 0 to 30 o Mo S o 70 e 90 jollo Cobito o All Poul Cam, 1900) 6 $1500 at Zero copy, 1900-IS M O

b & c)
copy costs
High 105000 6000
Low 35000 3200
difference 70000 2800
Variable cost = 2800/70000 = 0.04
Fixed cost = 3200 - (35000*0.04)=1800
d) Equation :
y = 0.04x + 1800
e) at 80000 copies, cost estimated = 80000*0.04 + 1800 = $5000
The estimate is different from April actual cost of $5100 at
80000 copies because it is an average of high & low and not
the actual average of all figures. It is calculated on the basis
of equation drawn on High/Low basis.